Propane burns completely with 150% excess air during a steady-flow combustion process. If both the reagents and the products are kept at 25 ° C and an atmosphere and the water in the products exists in liquid form, determine : to.

A.The real balanced equation

B. Heat transfer (kJ / kmol)

C. Adiabatic flame temperature (K)
Answer #1

prequired In produit, co, zbooknol, Ho. a) Caltz +50₂ - 3co +H₂O b Leo 100 knop of Propane from stoichiomely For I knol GH8 58 = 184503.705(T-298 Dhran for 100kond Dhan2-204300000 k] & Bhran & Ose 02-2043.0.0000 + 184503705 CT-298 5204300000 +5498210Now 8 = Arran tai - o -204300000 - 16320000 +184503.70 (T-298) => 0 = -220620000+184503.705(T-278) 275602 164.262184503.705Tfrom mole balance all the product flow rate is calculated. If outlet and inlet temperature is at 25 degree Celsius then total heat transfer is heat of reaction which is calculated from heat of formation. But if water is condensed then it releases some amount of heat. So on that basis total heat is calculated. Heat released is negative and heat absorb is positive. Adiabatic flame temperature is temperature of product when no amount of heat is transferring from reactor.